Two investments are made totaling $8800. For a certain year, these investments yield $1326 in simple interest. Part of the $8800 is invested at 16% and part at 14%. Find the amount invested at each rate.
Answer by mananth(16949)

x + y 8800 ------------------------1
14.00% x + 16.00% y = $1,326.00
Multiply by 100
14 x + 16 y = $132,600.00 --------2

Multiply (1) by -14
we get

-14 x -14 y = -123200.00

Add this to (2)

0 x 2 y = $9,400.00

divide by 2

y =$4,700.00 investment at 16.00%
Balance $4,100.00 investment at 14.00%